OpenStudy (anonymous):

Find the sum of this infinite geometric series. 100+60+36+...

OpenStudy (phi):

do you know the formula? see https://en.wikipedia.org/wiki/Geometric_series#Formula

OpenStudy (phi):

to use the formula you need to know the "r" which is called the "common ratio" take the first two terms, 100 and 60 100 * r = 60 divide both sides by 100 r= 60/100 r= 0.6 also, a is the first term. a is 100